# Self-Supervised Transformers as Iterative Solution Improvers for Constraint Satisfaction

### v1

@@ 4.3 Results
Nurse Scheduling.
For the Nurse Scheduling problem, ConsFormer matches OR-Tools in solving all tasks across both in-distribution and out-of-distribution instances within the 10-second timeout. This high accuracy is expected, given the large number of feasible solutions for each instance, as detailed in Appendix C. However, solving this problem neurally is non-trivial, as the model must learn to balance multiple constraints within a single step of solution refinement. ANYCSP was not evaluated on this dataset due to their difficulty dealing with the AllDifferent constraint. This result highlights ConsFormer’s potential to generalize to more complex problems with diverse constraint structures.
4.4 Ablations
@@
Effect of subset improvement.
Effect of variable information as positional encodings.
Table 4 and 5 show the performance of ConsFormer with different positional encodings. The value displayed indicates the percentage of in-distribution test instances solved by the model running 1k iterations
@@ 5 Conclusion
We introduced ConsFormer, a self-supervised Transformer for iteratively solving Constraint Satisfaction Problems. We showed that our method, trained to perform a single step of solution improvement, is able to generalize to harder out-of-distribution instances at test time, outperforming supervised and reinforcement learning approaches. Future work includes exploring neural-symbolic approaches incorporating ConsFormer such as those discussed in Section 3.4, other performance boosting techniques such as self-improvement techniques to augment the training data (Lee et al., 2025), as well as extending ConsFormer to more problems and more constraints, with the goal of devising a general continuous approximation for constraints.
@@ Appendix D Model Details
Sudoku Graph-coloring-5 Graph-coloring-10 Nurse Scheduling
Layer Count 7 4 7 7
Head Count 3 3 3 3
Embedding Size 128 128 128 126
Selection Probability $p$ 0.5 0.3 0.3 0.3

### v2

@@ 4.4 Ablations
Effect of variable information as positional encodings.
Table 5 and 6 show the performance of ConsFormer with different positional encodings. The value displayed indicates the percentage of in-distribution test instances solved by the model running $1,000$ iterations
@@
5 Limitations and Future Work
In its current form, ConsFormer assumes a fixed constraint structure, the effects of which on solution feasibility are implicitly learned during training via the loss function. ConsFormer does not receive explicit constraint representations as input. Constraints which are “parametric”, e.g., a SAT clause in which some variables may be negated and some not, cannot be handled with the architecture described herein. It is possible to address this limitation through explicit constraint representations in the input; this is an important direction for future work.
Other future work includes exploring neural-symbolic approaches incorporating ConsFormer such as those discussed in Section 3.4, other performance boosting techniques such as self-improvement to augment the training data (Lee et al., 2025), as well as extending ConsFormer to more problems and more constraints, with the goal of devising a general continuous approximation for constraints.
@@ 6 Conclusion
We introduced ConsFormer, a self-supervised Transformer for iteratively solving Constraint Satisfaction Problems. We showed that our method, trained to perform a single step of solution improvement, is able to generalize to harder out-of-distribution instances at test time, outperforming supervised and reinforcement learning approaches.
Acknowledgments
We thank the anonymous reviewers for their insightful feedback, which helped us identify key limitations and promising directions for future work. This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project). Elias B. Khalil acknowledges support from the SCALE AI Research Chair program.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix D Model Details
Sudoku Graph-coloring-5 Graph-coloring-10 Nurse Scheduling MAXCUT
Layer Count 7 4 7 7 4
Head Count 3 3 3 3 3
Embedding Size 128 128 128 126 128
Selection Probability $p$ 0.5 0.3 0.3 0.3 1.0
Appendix E Effects of Gumbel Softmax
We use Gumbel-Softmax in ConsFormer to enable differentiable sampling of discrete variables, aligning with the discrete nature of CSPs.
To better understand why Gumbel-Softmax improves performance, we investigate whether the benefit stems from the stochasticity introduced by Gumbel noise or simply from producing sharper output distributions. To isolate these factors, we compare against a softmax variant with temperature control:
$\text{Softmax}_{\tau}(z_{i})=\frac{\exp\left(z_{i}/\tau\right)}{\sum_{j}\exp\left(z_{j}/\tau\right)}$
This variant allows us to control the sharpness of the output distribution without introducing stochasticity. Table 9 presents results for Sudoku and Graph Coloring (percentage of instances solved), and MAXCUT (absolute gap to best known values).
Results indicate that softmax with temperature performs competitively or even better on smaller-scale problems such as Sudoku and Graph Coloring instances. However, for larger problems like MAXCUT with thousands of variables, Gumbel-Softmax clearly outperforms temperature-controlled softmax.
This suggests that while sharper distributions (e.g., from low temperature) are beneficial, the stochasticity allows the model to generalize better across larger problem instances. The randomness can promote diversity in intermediate solutions which may help the model escape local optima over multiple inference steps.
Table 9: Comparison of ConsFormer with and without Gumbel-Softmax
Method Gumbel-Softmax Softmax
$\tau=0.1$ $\tau=1$ $\tau=0.1$ $\tau=1$
Sudoku 100 100 100 100
Sudoku OOD 77.74 83.71 85.67 73.72
Graph-Coloring-5 V=50 78.16 76.91 77.33 74.91
Graph-Coloring-5 V=100 42.50 41.08 42.66 35.33
Graph-Coloring-10 V=100 52.60 53.25 53.66 53.0
Graph-Coloring-10 V=200 11.92 12.75 12.92 12.75
MAXCUT $|V|{=}800$ 24.44 102.89 123.11 126.56
MAXCUT $|V|{=}1K$ 18.22 44.0 58.33 56.89
MAXCUT $|V|{=}2K$ 47.0 119.33 123.67 135.11
MAXCUT $|V|{\geq}3K$ 155.88 187.0 287.38 305.25
Appendix F ConsFormer with Multi-Start
A simple extension to enhance our model is the implementation of multi-start. Instead of a single initial solution continuously refined, we maintain a pool of candidate solutions that are updated concurrently. A solution is accepted as soon as any candidate satisfies all constraints. This approach naturally complements ConsFormer, as the Transformer architecture efficiently handles batched processing.
Table 10 reports results for various candidate pool sizes. A candidate count of 1 corresponds to the standard version of ConsFormer used in the main paper.
We observe that as the number of candidates increases, the number of update iterations per candidate slightly decreases due to the fixed time budget. However, the drop is modest compared to the increase in candidates, highlighting the scalability of the Transformer-based model. We observe that the multi-start strategy is able to noticably improve model performance, boosting accuracy from 47.33% to 55.67% for graph coloring with $k=5$ and 11.92% to 15.00% for $k=10$.
These findings highlight the potential of integrating with symbolic strategies, such as restarts and backtracking, as discussed in Section 3.4. We leave further exploration of these hybrid techniques to future work.
Table 10: Performance comparison for Graph-Coloring tasks on Out-of-Distribution evaluation for ConsFormer. Candidates Count refers to the number of solutions used for multi-start. # Iterations Average shows the number of iterations each candidate went through under the time limit.
Method Harder OOD Pool # Iterations
Instances Size Avg
Graph-Coloring-5 ($n=100$)
OR-Tools (10s) 57.16 - -
ConsFormer (10s) 47.33 1 2310
ConsFormer (10s) 43.42 2 2213
ConsFormer (10s) 50.92 10 1634
ConsFormer (10s) 55.17 50 1613
ConsFormer (10s) 55.67 100 892
Graph-Coloring-10 ($n=200$)
OR-Tools (10s) 10.25 - -
ConsFormer (10s) 11.92 1 1490
ConsFormer (10s) 13.67 2 1445
ConsFormer (10s) 13.92 5 1064
ConsFormer (10s) 14.42 10 1150
ConsFormer (10s) 15.00 50 806
ConsFormer (10s) 13.25 100 229
Appendix G Additional Baselines for Graph Coloring
We provide additional baselines for comparison on the Graph Coloring task. We first run OR-Tools for additional time (30 and 60 seconds) with 10 colors (where ConsFormer had previously outperformed it under 10s). We see that CP-SAT can outperform our method on small instances (nodes=100) with extended time, but it still underperforms on larger instances (nodes=200), even with 6x more time. Furthermore, in the new MAXCUT problem, 20 parallel runs–each with 180s limit–were used to compute the results. ConsFormer also outperforms OR-Tools by a significant margin.
We then include 3 additional heuristic baselines: Greedy Coloring, Feasibility Jump, Random Search. The first is the greedy coloring algorithm implemented by networkx and the other two are local search approaches implemented by OR-Tools. Results are shown in Table 11. We observe that while the local-search based heuristics were able to perform well on the smaller instances, their performance significantly worsens on the larger instances with 10 colors.
Method Test Harder OOD
Instances Instances
Graph-Coloring-5 ($n=50\rightarrow n=100$)
Greedy 32.42 0.0
OR-Tools-FJ (10s) 82.83 54.5
OR-Tools-RS (10s) 83.08 56.91
OR-Tools (10s) 83.08 57.16
ANYCSP (10s) 79.17 34.83
ConsFormer (10s) 81.00 47.33
Graph-Coloring-10 ($n=100\rightarrow n=200$)
Greedy 0.75 0.0
OR-Tools-FJ (10s) 35.66 6.0
OR-Tools-RS (10s) 49.75 9.08
OR-Tools (10s) 52.41 10.25
OR-Tools (30s) 53.58 11.16
OR-Tools (60s) 53.67 11.66
ANYCSP (10s) 0.00 0.00
ConsFormer (10s) 52.60 11.92
Appendix H Penalty Functions Design
As detailed in Section 3.3, our continuous penalties are designed such that
$p(X)=0\iff c(X)=\texttt{True},$
where penalty $p$ approximates constraint $c$ defined over variables $X$. Intuitively, the penalty $p$ only evaluates to 0 when the constraint $c$ is satisfied.
This approach follows constraint-based local search (Hentenryck & Michel, 2009). Here, a constraint $c$ is associated with a “violation degree” function $v_{c}$ where
$v_{c}(X)=0\iff c(X)=\texttt{satisfied}$
Specific functions to evaluate violation degrees are designed for different global constraints. For example, a violation function for $\texttt{AllDifferent}(x_{1},\ldots,x_{n})$ can be defined as
$v_{c}(x_{1},\ldots,x_{n})=\sum_{i\in S}\max\left(0,\ |\{x_{j}=i\mid j\in 1,\ldots,n\}|-1\right)$
where $S$ is the set of all values in the domains of $x_{1},\ldots,x_{n}$. Intuitively, this violation degree counts how many values are assigned to more than one variable among $x_{1},\ldots,x_{n}$. This idea is extended to design the continuous penalty function for $\textsc{AllDifferent}_{m>n}(x_{1},\ldots,x_{n})$.
The design of the penalty functions is a flexible and modular component of our framework, and can benefit from further improvements which we leave for future work.
Appendix I Direct Gradient Descent on Variables
A natural baseline to consider is optimizing variable assignments directly using stochastic gradient descent (SGD), without a learned architecture. Specifically, if we ignore the Transformer component of ConsFormer and instead treat the variable assignments as continuous parameters, we can optimize them using our self-supervised loss. In theory, this procedure should lead to a relaxed satisfying solution.
However, we found that in practice, this method frequently converges to poor local optima. Additionally, because the optimization is performed independently for each instance, the updates cannot generalize to other instances.
To illustrate the limitation of this approach, we performed a simple experiment on Sudoku. Starting from a random initialization of missing cells, we applied SGD for 10000 steps using the self-supervised loss. The table below reports the average number of satisfied AllDifferent constraints (out of 27 total) across 10 runs:
Table 12: Direct SGD optimization of Sudoku variable assignments
# Missing Cells 19 33 41 47
# Satisfied AllDifferent 26.8 25.8 24.5 21.8
As expected, performance degrades as the number of missing cells increases. The optimization becomes harder, and the model fails to satisfy all constraints. This highlights the importance of learning a iterative improvement model, rather than relying solely on instance-specific gradient descent.